# Effects of anisotropy in an anisotropic extension of $w$CDM model

### v1

@@ IV Results and Discussion
In Table 3, we summarize the observational constraints on free parameters and some derived parameters of the $w$CDM and $w$CDM+$\Omega_{\rm\sigma 0}$ models at $68\%$CL from two data combinations: BAO+CC+BBN+PP and BAO+CC+BBN+PPSH0ES. The second row over each parameter (mentioned in blue color) represents the constraints at $68\%$CL on $\Lambda$CDM and $\Lambda$CDM+$\Omega_{\rm\sigma 0}$ models from same set of data combinations. The upper bounds on anisoptropic expansion parameter, $\Omega_{\rm\sigma 0}$, at $95\%$ CL are of order $10^{-13}$ for $w$CDM+$\Omega_{\rm\sigma 0}$ model with both the data combinations. The upper bounds of $\Omega_{\rm\sigma 0}$ for this model read as: $\Omega_{\rm\sigma 0}<2.3\times 10^{-13}$ and $\Omega_{\rm\sigma 0}<4\times 10^{-13}$ from BAO+CC+BBN+PP and BAO+CC+BBN+PPSH0ES, respectively. On other hand, the upper bounds on anisoptropic expansion parameter $\Omega_{\rm\sigma 0}$ for $\Lambda$CDM+$\Omega_{\rm\sigma 0}$ model are: $\Omega_{\rm\sigma 0}<2.0\times 10^{-14}$ and $\Omega_{\rm\sigma 0}<6.5\times 10^{-14}$ from BAO+CC+BBN+PP and BAO+CC+BBN+PPSH0ES combinations, respectively. Thus, we see that the $w$CDM+$\Omega_{\rm\sigma 0}$ model provides weaker upper bounds on anisotropy parameter by one order of magnitude as compared to $\Lambda$CDM+$\Omega_{\rm\sigma 0}$ model with both the data combinations. We observe that anisotropic parameter $\Omega_{\rm\sigma 0}$ has a strong positive correlation with DE EoS parameter from both the data combinations. This means that larger amount of anisotropy in the Universe results larger value of DE EoS, see Figure 1.
Table 3: Constraints at 68% CL on the free and some derived parameters of the $w$CDM and $w$CDM+$\Omega_{\rm\sigma 0}$ models from BAO+CC+BBN+PP and BAO+CC+BBN+PPSH0ES data. The upper bounds on $\Omega_{\rm\sigma 0}$ at 95% CL are displayed. The Hubble constant $H_{0}$ is measured in the unit of $\rm km\,s^{-1}\rm Mpc^{-1}$. The entries in second row over each parameter (in blue colour) represent the constraints on $\Lambda$CDM and $\Lambda$CDM+$\Omega_{\rm\sigma 0}$ model parameters.
Data set BAO+CC+BBN+PP BAO+CC+BBN+PP+SH0ES
@@
Figure 1: Two-dimensional marginalized confidence regions at 68% and 95% CL for $\Omega_{\rm\sigma 0}$ and $w_{\rm de0}$ for $w$CDM+$\Omega_{\rm\sigma 0}$ model from BAO+CC+BBN+PP and BAO+CC+BBN+PPSH0ES data combinations.
@@
For the $w$CDM+$\Omega_{\rm\sigma 0}$ model, it can be observed from Table 3 that the DE EoS parameter does not prefer the cosmological constant form of DE (having $w_{\rm de0}=-1$) with both data combinations. Also, at $95\%$ CL, the DE EoS parameter ($w_{\rm de0}=-0.853^{+0.096}_{-0.100}$ for BAO+CC+BBN+PP data combination and $w_{\rm de0}=-0.88^{+0.10}_{-0.11}$ for BAO+CC+BBN+PPSH0ES data combination) does not prefer the cosmological constant form of DE with both data combinations. For this model, the mean values of DE EoS parameter (at $68\%$ CL as well as $95\%$ CL range), $w_{\rm de0}$, lie in quintessence region $(w_{\rm de0}>-1)$, pointing towards quintessence behaviour of DE with both data combinations. Also, for the $w$CDM model, the mean value of $w_{\rm de0}$ at $68\%$ CL, as shown in Table 3, lies in quintessence region with BAO+CC+BBN+PP data combination whereas, it touches the phantom barrier at $95\%$ CL with the same data combination ($w_{\rm de0}=-0.930^{+0.086}_{-0.090}$ for BAO+CC+BBN+PP data combination). Again, for the same model with BAO+CC+BBN+PPSH0ES data combination, the mean value of $w_{\rm de0}$ lies in phantom region $(w_{\rm de0}<-1)$ at $68\%$ CL as well as $95\%$ CL. At $95\%$ CL, $w_{\rm de0}=-1.062^{+0.078}_{-0.082}$ for BAO+CC+BBN+PPSH0ES data combination. Also, we observe from Figure 2 that $w_{\rm de0}$ finds a strong negative correlation with $H_{0}$ in case of $w$CDM model whereas slight negative correlation in case of $w$CDM+$\Omega_{\rm\sigma 0}$ model from both the data combinations.
@@
Figure 2: Two-dimensional marginalized confidence regions at 68% and 95% CL of $H_{0}$ and $w_{\rm de0}$ for $w$CDM and $w$CDM+$\Omega_{\rm\sigma 0}$ models from BAO+CC+BBN+PP and BAO+CC+BBN+PPSH0ES data combinations. The vertical yellow band represents $H_{\rm 0}^{\rm R22}=73.04\pm 1.04\leavevmode\nobreak\ {\rm km\,s^{-1}\,Mpc^{-1}}$ from SH0ES measurement. The horizontal black line represents $w_{\rm de0}=-1$.
@@
Figure 3: One-dimensional marginalized distributions of $\Omega_{\rm m0}$ for $\Lambda$CDM, $\Lambda$CDM+$\Omega_{\rm\sigma 0}$, $w$CDM and $w$CDM+$\Omega_{\rm\sigma 0}$ models from BAO+CC+BBN+PP (left panel) and BAO+CC+BBN+PPSH0ES (right panel) data combinations.
One of the most well known and intriguing puzzle nowadays in cosmology is discrepancy between Planck CMB data within the baseline $\Lambda$CDM model and direct local measurements of the current rate of expansion of the Universe (that is, the value of Hubble constant, $H_{0}$), as discussed in Section I. Now, we discuss the constraints on Hubble constant for the considered models and compare the results with the previous work [60]. For BAO+CC+BBN+PP data combination, the constraint on $H_{0}$ for $w$CDM ($\Lambda$CDM) model reads: $H_{\rm 0}=66.4\pm 1.3\,(68.02\pm 0.84)\leavevmode\nobreak\ {\rm km\,s^{-1}\,Mpc^{-1}}$, and for $w$CDM+$\Omega_{\rm\sigma 0}$ ($\Lambda$CDM+$\Omega_{\rm\sigma 0}$) model, it reads: $H_{\rm 0}=68.4^{+1.6}_{-1.6}\,(70.1^{+1.2}_{-1.5})\leavevmode\nobreak\ {\rm km\,s^{-1}\,Mpc^{-1}}$ at 68% CL. For BAO+CC+BBN+PP data combination, we have obtained lower mean values of $H_{0}$ for $w$CDM and $w$CDM+$\Omega_{\rm\sigma 0}$ models as compared to $\Lambda$CDM and $\Lambda$CDM+$\Omega_{\rm\sigma 0}$ models, respectively. For BAO+CC+BBN+PPSH0ES data combination, the constraint on $H_{0}$ for $w$CDM ($\Lambda$CDM) model reads: $H_{\rm 0}=71.51\pm 0.83\,(70.79\pm 0.69)\leavevmode\nobreak\ {\rm km\,s^{-1}\,Mpc^{-1}}$, and for $w$CDM+$\Omega_{\rm\sigma 0}$ ($\Lambda$CDM+$\Omega_{\rm\sigma 0}$) model, it reads: $H_{\rm 0}=72.24\pm 0.82\,(72.67\pm 0.85)\leavevmode\nobreak\ {\rm km\,s^{-1}\,Mpc^{-1}}$ at 68% CL. Thus, from BAO+CC+BBN+PPSH0ES data $w$CDM model provides higher mean value of $H_{0}$ as compared to $\Lambda$CDM model, whereas $w$CDM+$\Omega_{\rm\sigma 0}$ and $\Lambda$CDM+$\Omega_{\rm\sigma 0}$ models provide more or less the same mean values of $H_{0}$. The highest value of Hubble constant $H_{0}$ obtained in the present analysis is $72.24\pm 0.82\leavevmode\nobreak\ {\rm km\,s^{-1}\,Mpc^{-1}}$ from BAO+CC+BBN+PPSH0ES data combinations which is consistent with $H_{0}^{\rm R22}=73.04\pm 1.04\leavevmode\nobreak\ {\rm km\,s^{-1}\,Mpc^{-1}}$ from SH0ES measurement. Quantifying the $H_{0}$ tension with SH0ES measurement for BAO+CC+BBN+PP data combination, we found that there is $3.9\sigma$ tension on $H_{0}$ in the $w$CDM model and there is $2.4\sigma$ tension on $H_{0}$ in $w$CDM+$\Omega_{\rm\sigma 0}$ model. Also, for BAO+CC+BBN+PPSH0ES data combination, $1.1\sigma$ tension on $H_{0}$ is observed in $w$CDM model and $0.6\sigma$ tension is observed on $H_{0}$ in $w$CDM+$\Omega_{\rm\sigma 0}$ model. So, for BAO+CC+BBN+PP data combination $H_{0}$ tension is relieved by $1.5\sigma$ in presence of anisotropy of the order $10^{-13}$ with constant EoS parameter of DE. Also, for BAO+CC+BBN+PPSH0ES data combination $H_{0}$ tension is relieved by $0.5\sigma$ in presence of anisotropy of the order $10^{-13}$ with constant EoS parameter of DE. See the Figure 2 where vertical yellow band represents $H_{\rm 0}^{\rm R22}=73.04\pm 1.04\leavevmode\nobreak\ {\rm km\,s^{-1}\,Mpc^{-1}}$ from SH0ES measurement and constraints on $H_{0}$ in Table 3. Also, from Figure 4, we observe a positive correlation of $H_{0}$ with $\Omega_{\rm\sigma 0}$ with both the data combinations. So, in order to get a larger value of $H_{0}$, a larger amount of anisotropy is required.
@@
Figure 4: One and two dimensional marginalized confidence regions at 68% and 95% CL for some selected parameters of $w$CDM and $w$CDM+$\Omega_{\rm\sigma 0}$ models from BAO+CC+BBN+PP and BAO+CC+BBN+PPSH0ES data combinations.
@@
Now, we discuss how $w_{\rm de0}$ and $\Omega_{\rm\sigma 0}$ affect the present day matter density parameter, $\Omega_{\rm m0}$. Firstly, we observe that the $w$CDM model provides significantly lower mean values of $\Omega_{\rm m0}$ as compared to $\Lambda$CDM model for both data combinations. Second, adding anisotropy to $\Lambda$CDM model slightly reduces the mean value of $\Omega_{\rm m0}$ for both the data combinations.Also, inclusion of anisotropy into the $w$CDM model significantly reduces the mean values of $\Omega_{\rm m0}$ for both the data combinations. The constraints on $\Omega_{\rm m0}$ for $w$CDM model are: $\Omega_{\rm m0}=0.258\pm 0.015$ and $\Omega_{\rm m0}=0.287\pm 0.012$ from the data combinations BAO+CC+BBN+PP and BAO+CC+BBN+PPSH0ES, respectively. The constraints on $\Omega_{\rm m0}$ for $w$CDM+$\Omega_{\rm\sigma 0}$ model are: $\Omega_{\rm m0}=0.230\pm 0.018$ for the data combination BAO+CC+BBN+PP and $\Omega_{\rm m0}=0.233\pm 0.018$ for the data combination BAO+CC+BBN+PPSH0ES. We can observe the deviations in the distribution of $\Omega_{\rm m0}$ for considered models in one dimensional marginalized probability distribution as shown in Figure 3 (also, see table 3) for both the data combinations. We also observe that parameters, $\Omega_{\rm\sigma 0}$ and $\Omega_{\rm m0}$ finds a negative correlation with both data combinations, see $\Omega_{\rm\sigma 0}-\Omega_{\rm m0}$ parametric space in the Figure 4.
@@
The considered $w$CDM and $w$CDM+$\Omega_{\rm\sigma 0}$ models provides higher values of $r_{\rm d}$ as compared to $\Lambda$CDM and $\Lambda$CDM+$\Omega_{\rm\sigma 0}$ models with BAO+CC+BBN+PP data combination. Also, from BAO+CC+BBN+PPSH0ES data combination, we obtain the result that the $w$CDM model prefers a lower value of $r_{\rm d}$ than $\Lambda$CDM model, but $\Lambda$CDM+$\Omega_{\rm\sigma 0}$ and $w$CDM+$\Omega_{\rm\sigma 0}$ provide equivalent mean values on $r_{\rm d}$ with similar error bars. Further, we notice that the addition of anisotropy reduces the value of $r_{\rm d}$ in both the cases. We can observe from Figure 4, that the parameters $\Omega_{\rm\sigma 0}$ and $r_{\rm d}$ are negatively correlated with both the data combinations, which means that for a small increment in $\Omega_{\rm\sigma 0}$ values, there would be corresponding decrement in $r_{\rm d}$ values.
@@
We have observed that different model parameters are influenced due to anisotropy. For instance, we observe a negative correlation between $H_{0}$ and $w_{\rm de0}$ which can be seen in Figure 2. Note that parameters influence each other while fitting the multi-parameter space simultaneously [81, 82]. Also, the other parameters of the models under consideration (with constant EoS of DE) are affected by anisotropy in such a way that results into the relieving of the $H_{0}$ tension up to $\sim 1\sigma$. It is important to emphasize that in equation (19), the scaling $a^{-6}$ has relevance for the early Universe while it becomes negligible in the late Universe. Thus, the anisotropic model under consideration behaves like the Early DE model. For more about Early DE models, we refer the readers to see the recent and interesting review [33] and references therein.

### v2

@@ IV Results and Discussion
In Table 3, we summarize the constraints on free parameters and some derived parameters of the $w$CDM and $w$CDM+$\Omega_{\rm\sigma 0}$ models at $68\%$CL from both combinations: $D_{1}$ and $D_{2}$. The second row over each parameter (mentioned in blue color) represents the constraints at $68\%$CL for $\Lambda$CDM and $\Lambda$CDM+$\Omega_{\rm\sigma 0}$ models from both data sets. The upper bounds on the anisotropic expansion parameter, $\Omega_{\rm\sigma 0}$ at $95\%$ CL are of order $10^{-13}$ for $w$CDM+$\Omega_{\rm\sigma 0}$ model with both $D_{1}$ and $D_{2}$. The upper bounds of $\Omega_{\rm\sigma 0}$ for this model read as: $\Omega_{\rm\sigma 0}<2.3\times 10^{-13}$ and $\Omega_{\rm\sigma 0}<4\times 10^{-13}$ from $D_{1}$ and $D_{2}$, respectively. On the other hand, the upper bounds on the anisotropic expansion parameter $\Omega_{\rm\sigma 0}$ for $\Lambda$CDM+$\Omega_{\rm\sigma 0}$ model are: $\Omega_{\rm\sigma 0}<2.0\times 10^{-14}$ and $\Omega_{\rm\sigma 0}<6.5\times 10^{-14}$ from $D_{1}$ and $D_{2}$, respectively. Thus, we see that the $w$CDM+$\Omega_{\rm\sigma 0}$ model provides weaker upper bounds on the anisotropy parameter by one order of magnitude as compared to $\Lambda$CDM+$\Omega_{\rm\sigma 0}$ model with both the data combinations. We observed that the anisotropic parameter $\Omega_{\rm\sigma 0}$ finds a strong positive correlation with the DE EoS parameter from $D_{1}$ and $D_{2}$. This means that a larger amount of anisotropy in the Universe results in a larger value of the DE EoS, see Figure 1.
Table 3: Constraints at 68% CL on the free and some derived parameters of the $w$CDM and $w$CDM+$\Omega_{\rm\sigma 0}$ models from $D_{1}$ and $D_{2}$ data. The upper bounds on $\Omega_{\rm\sigma 0}$ are displayed at 95% CL. The Hubble constant $H_{0}$ is measured in the unit of $\rm km\,s^{-1}\rm Mpc^{-1}$. The entries in the second row over each parameter (in blue color) represent the constraints on $\Lambda$CDM and $\Lambda$CDM+$\Omega_{\rm\sigma 0}$ model parameters.
Data set $D_{1}$ $D_{2}$
@@
Figure 1: Two-dimensional marginalized confidence regions at 68% and 95% CL for $\Omega_{\rm\sigma 0}$ and $w_{\rm de0}$ for the $w$CDM+$\Omega_{\rm\sigma 0}$ model from $D_{1}$ and $D_{2}$ data.
@@
For the $w$CDM+$\Omega_{\rm\sigma 0}$ model, it can be observed from Table 3 that the DE EoS parameter does not prefer the cosmological constant form of the DE (having $w_{\rm de0}=-1$) with $D_{1}$ and $D_{2}$. Also, at $95\%$ CL, the DE EoS parameter ($w_{\rm de0}=-0.853^{+0.096}_{-0.100}$ for $D_{1}$ and $w_{\rm de0}=-0.88^{+0.10}_{-0.11}$ for $D_{2}$) does not prefer the cosmological constant form of the DE. For this model, the mean values of the DE EoS parameter (at $68\%$ CL as well as $95\%$ CL) lie in the quintessence region $(w_{\rm de0}>-1)$, pointing towards the quintessence behavior of the DE. Also, for the $w$CDM model, the mean value of $w_{\rm de0}$ at $68\%$ CL, as shown in Table 3, lies in the quintessence region with $D_{1}$ whereas, it touches the phantom barrier at $95\%$ CL with $D_{1}$ ($w_{\rm de0}=-0.930^{+0.086}_{-0.090}$). Again, for the same model with $D_{2}$, the mean value of $w_{\rm de0}$ lies in the phantom region $(w_{\rm de0}<-1)$ at $68\%$ CL as well as $95\%$ CL with $w_{\rm de0}=-1.062^{+0.078}_{-0.082}$ at $95\%$ CL. Also, we observe from Figure 2 that $w_{\rm de0}$ finds a strong negative correlation with $H_{0}$ in the case of $w$CDM model whereas a slight negative correlation in the case of $w$CDM+$\Omega_{\rm\sigma 0}$ model with both $D_{1}$ and $D_{2}$ combinations.
@@
Figure 2: Two-dimensional marginalized confidence regions at 68% and 95% CL of $H_{0}$ and $w_{\rm de0}$ for the $w$CDM and $w$CDM+$\Omega_{\rm\sigma 0}$ models from $D_{1}$ and $D_{2}$ data. The vertical yellow band represents $H_{\rm 0}^{\rm R22}=73.04\pm 1.04\leavevmode\nobreak\ {\rm km\,s^{-1}\,Mpc^{-1}}$ from SH0ES measurement. The horizontal dotted black line represents $w_{\rm de0}=-1$.
[…]
@@
Figure 4: One-dimensional marginalized probability distributions of $\Omega_{\rm m0}$ for the considered models from $D_{1}$ (left) and $D_{2}$ (right) combinations.
[…]
@@
Figure 5: One and two dimensional marginalized confidence regions at 68% and 95% CL for some selected parameters of the $w$CDM and $w$CDM+$\Omega_{\rm\sigma 0}$ models from $D_{1}$ and $D_{2}$ combinations.
Figure 6: One-dimensional marginalized probability distributions of $H_{0}$ for the $w$CDM and $w$CDM+$\Omega_{\rm\sigma 0}$ models from $D_{1}$ and PPSH0ES data.
@@
Now, we discuss how the present DE EoS, $w_{\rm de0}$ and $\Omega_{\rm\sigma 0}$ affect the present day matter density parameter, $\Omega_{\rm m0}$. Firstly, we observed that the $w$CDM model provides significantly lower mean values of $\Omega_{\rm m0}$ compared to the $\Lambda$CDM model from $D_{1}$ and $D_{2}$. Second, adding anisotropy to the $\Lambda$CDM model slightly reduces the mean value of $\Omega_{\rm m0}$ for both combinations. Also, the inclusion of anisotropy into the $w$CDM model significantly reduces the mean values of $\Omega_{\rm m0}$ for $D_{1}$ and $D_{2}$. The constraints on $\Omega_{\rm m0}$ for the $w$CDM model are: $\Omega_{\rm m0}=0.258\pm 0.015$ and $\Omega_{\rm m0}=0.287\pm 0.012$ from $D_{1}$ and $D_{2}$, respectively. The constraints on $\Omega_{\rm m0}$ for the $w$CDM+$\Omega_{\rm\sigma 0}$ model are: $\Omega_{\rm m0}=0.230\pm 0.018$ from $D_{1}$ and $\Omega_{\rm m0}=0.233\pm 0.018$ from $D_{2}$. We can observe the deviations in the distribution of $\Omega_{\rm m0}$ for considered models in one-dimensional marginalized probability distribution as shown in Figure 4 (also, see table 3). We also observe that the parameters, $\Omega_{\rm\sigma 0}$ and $\Omega_{\rm m0}$ finds a negative correlation with both data combinations, see $\Omega_{\rm\sigma 0}-\Omega_{\rm m0}$ parametric space in Figure 5.
@@
The considered $w$CDM and $w$CDM+$\Omega_{\rm\sigma 0}$ models provides higher values of $r_{\rm d}$ as compared to the $\Lambda$CDM and $\Lambda$CDM+$\Omega_{\rm\sigma 0}$ models with $D_{1}$. Also, from $D_{2}$, we obtain the result that the $w$CDM model prefers a lower value of $r_{\rm d}$ than the $\Lambda$CDM model, but the $\Lambda$CDM+$\Omega_{\rm\sigma 0}$ and the $w$CDM+$\Omega_{\rm\sigma 0}$ provide equivalent mean values on $r_{\rm d}$ with similar error bars. Further, we notice that adding the anisotropy reduces the value of $r_{\rm d}$ in both cases. We can observe from Figure 5, that the parameters $\Omega_{\rm\sigma 0}$ and $r_{\rm d}$ are negatively correlated, which means that for a small increment in $\Omega_{\rm\sigma 0}$ values, there would be corresponding decrement in $r_{\rm d}$ values.
@@
We have observed that different model parameters are influenced by anisotropy. For instance, we observe a negative correlation between $H_{0}$ and $w_{\rm de0}$ which can be seen in Figure 2. Note that parameters influence each other while fitting the multi-parameter space simultaneously [82, 83]. Also, the other parameters of the models under consideration (with constant EoS of DE) are affected by anisotropy in such a way that relieves the $H_{0}$ tension up to $\sim 1\sigma$. It is important to emphasize that in equation (19), the scaling $a^{-6}$ has relevance for the early Universe while it becomes negligible in the late Universe. Thus, the anisotropic model under consideration behaves like the Early DE model. For more about Early DE models, we refer the readers to the recent and interesting review [33] and references therein.
IV.1 STATISTICAL MODEL COMPARISON
The model comparison criteria, play a significant role in evaluating the sustainability of the models. We apply well-known statistical criteria here: the Akaike Information Criteria (AIC) [84, 85]. The AIC is used to figure out the goodness-of-fit of different models. The model with the lowest AIC is the most preferred by observations. The AIC considers the maximum likelihood function and the number of free parameters of the model. It is defined as
$\text{AIC}=-2\ln\mathcal{L}_{\rm max}+2\,\texttt{N}\quad=\chi_{\rm min}^{2}+2\,\texttt{N},$ (34)
where $\mathcal{L}_{\rm max}$ indicates the maximum likelihood function and N denotes the total number of free parameters in the model. To compare the derived model with a reference model, the difference of AIC values, i.e., $\Delta\text{AIC}=\text{AIC}_{\rm model}-\text{AIC}_{\rm reference}$ play a crucial role than the AIC values itself. One may say a model is more favored than the other if $\Delta\text{AIC}$ is greater than a threshold value, $\Delta_{\rm threshold}$ [86]. The thumb rule of AIC states that a universal threshold, $\Delta_{\rm threshold}=5$ can be used for the model comparison. In [87], the author mentioned that depending on the minimum threshold AIC difference, one model is preferred over the other.
Table 4: The difference, $\Delta\text{AIC}=\text{AIC}_{\rm model}-\text{AIC}_{\rm\Lambda CDM}$ of considered models with respect to $\Lambda$CDM model with $D_{1}$ and $D_{2}$.
Data set $\Lambda$CDM+ $\Omega_{\sigma 0}$ $w$CDM $w$CDM + $\Omega_{\sigma 0}$
$D_{1}$ $2.38$ 00 $-0.12$
$D_{2}$ $-9.5$ $-0.3$ $-10.5$
The $\Delta\rm AIC$ values of the considered models (with respect to the $\Lambda$CDM) are shown in Table 4 for $D_{1}$ and $D_{2}$ combinations. We see that from $D_{1}$, the difference between the models is not statistically significant as the values are less than $\Delta_{\rm threshold}$. From $D_{2}$, we observe that the $w$CDM model has no statistical preference over the standard model whereas the $\Lambda$CDM+ $\Omega_{\sigma 0}$ and $w$CDM + $\Omega_{\sigma 0}$ show a strong preference over the $\Lambda$CDM model. Observing a preference via the Bayesian model comparison is not surprising as we know that the last term in the AIC acts as a penalty for models with many parameters.
In Figure 6, we show one-dimensional marginalized probability distributions of $H_{0}$ for the $w$CDM and $w$CDM+$\Omega_{\rm\sigma 0}$ models from $D_{1}$ and PPSH0ES data. We notice a significant overlapping of $H_{0}$ values within $w$CDM+$\Omega_{\rm\sigma 0}$ model with $D_{1}$ and PPSH0ES data but that is not the case within the $w$CDM model. So it may be treated as a statistical caveat of the joint analyses with the PPSH0ES data that there exist significant tensions among $D1$ and PPSH0ES datasets within the models under consideration, especially within the $w$CDM model.